please help me with this problem.
r + (-1\2)= -3\4

Answer by nerdybill(7384)

r + (-1\2)= -3\4
.
I'm assuming your "\" is division.
If so, start by multiplying both sides by 4 to get:
4r + (-2)= -3
4r - 2 = -3
Now, add 2 to both sides to get:
4r = -1
Finally divide both sides by 4:
r = -1/4